### v1

@@ 3.2 The Reduction Technique
    There exists a $B\in\mathbf{\Omega}^{n\times n}_{[d]}$ such that $h(x)=B(x)x$.
@@
Proof 3.12.
First, we establish (13). Consider a linear change of variable $y=Tx$, for $T\in\mathbb{R}^{n\times n}$ nonsingular and $y_{1}:=x^{\top}\mathbf{1}$ (i.e., the first row of $T$ is $\mathbf{1}^{\top}$). Then, $x^{\top}g(x)$ can be written as a polynomial in $y$, which we denote by $p(y)$. Since $x^{\top}g(x)=0$ for all $x\in\mathbf{H}$, we have that $p(y)=0$ for all $y$ with $y_{1}=1$. It follows that $(1-y_{1})$ is a factor of $p(y)$ and hence, $(1-x^{\top}\mathbf{1})$ is a factor of $x^{\top}g(x)$.
@@
Proof 3.14.
Since $s$ does not contain a constant or a linear term, we can write
@@
Proof 3.16.
We establish the proposition by a direct application of Lemma 3.13. For any $s(x)\in\mathbf{P}_{[d+2]}^{1}$, the following is true
$(1-x^{\top}\mathbf{1})s(x)$ $=x^{\top}H(x)x(1-x^{\top}\mathbf{1})$
$=x^{\top}(H(x)x-x^{\top}H(x)x\mathbf{1})$
Denote $g(x)=H(x)x-x^{\top}H(x)x\mathbf{1}$, we have
$x^{\top}g(x)=\left(1-x^{\top}\mathbf{1}\right)s(x),\text{ for any }x\in\mathbb{R}^{n},$
holds, and this completes the proof. ∎
@@
where the third equality follows from the fact that $\mathbf{1}x=1$ for $x\in\mathbf{H}$ and the last equality follows from the definition (14) of $h$. ∎
@@
As a consequence, if $H(x)$ is the payoff matrix (so $H(x)x$ is the payoff vector [14]), then the associated replicator dynamics are given by
@@
where $g\in\mathbf{P}_{[2]}^{n}$, and $x^{\top}g(x)=\left(1-x_{1}-x_{2}-x_{3}\right)x_{1}^{2}+2x_{1}^{2}x_{2}-2x_{1}^{2}x_{2}=0$, for any $x\in\mathbf{H}$. We claim that there does not exists an $H\in\mathbf{P}_{[0]}^{n\times n}$ such that $g(x)$ can be written in the form of (16). Suppose, to the contrary, that such constant matrix $H$ exists; then, it follows that for any $i,j\in\{1,2,3\}$, the difference $g_{i}(x)-g_{j}(x)=(Hx)_{i}-(Hx)_{j}$ is a linear function in $x$, which contradicts the fact that $g_{2}(x)-g_{3}(x)=-x_{1}^{2}$ is quadratic. We thus conclude that no such constant $H$ exists for (16) to hold.
@@ 3.3 Proof of Theorem 1
The procedure of constructing the skew-symmetric $A(x)$ such that $g(x)=A(x)x$ is summarized in Algorithm 2.
@@ 8 Input $g^{\prime}(x)$ for Algorithm 1 to obtain $A^{\prime}(x)$
Output: Skew-symmertic payoff matrix $A(x)=A^{\prime}(x)+B(x)$
@@ 4 Conclusion and Outlook
Future work can extend the theory beyond polynomial payoffs to analytic or otherwise smooth maps, broadening its scope. Moreover, the conditions under which the construction in Example 3.19 can be realized remain to be explored. By combining the present results with local analysis of replicator dynamics, one can translate the behavior of systems driven by higher-degree payoff matrices into equivalent, lower-degree formulations, gaining insight into otherwise intractable systems.
@@ Appendix A Nullspace of map $\phi$
Given $H\in\mathbf{P}_{[0]}^{n\times n}$, and $\phi$ as defined in (2), the nullspace of $\phi$ is $\left\{\mathbf{1}v^{\top}\mid v\in\mathbb{R}^{n}\right\}$.
Proof A.21.
@@
Sufficiency. Let $H=\mathbf{1}v^{\top}$ with some $v\in\mathbb{R}^{n}$. For any $x\in\mathbb{R}^{n}$ with $x^{\top}\mathbf{1}=1$, we have
@@
$\phi(x)=Hx-\left(x^{\top}Hx\right)\mathbf{1}=\mathbf{1}\left(v^{\top}x\right)-\mathbf{1}\left(v^{\top}x\right)=0.$
@@
Necessity. Conversely, suppose $\phi(H)=0$ for all $x$ with $x^{\top}\mathbf{1}=1$. Set $x=e_{i}$, then we have
$Hx=He_{i},x^{\top}Hx=e_{i}^{\top}He_{i}=h_{ii},$
and the following must hold for any $i\in\{1,2,\ldots,n\}$:
$\phi(x)=\operatorname{diag}(x)(Hx-x^{\top}Hx\mathbf{1})=\operatorname{diag}(e_{i})(He_{i}-h_{ii}\mathbf{1})=0,$
which implies that the $i$th column of $H$ is $h_{ii}\mathbf{1}$. Moreover, $H=\mathbf{1}[h_{11},h_{22},\ldots,h_{nn}]$. Set $v^{\top}=[h_{11},h_{22},\ldots,h_{nn}]$, we have $H=\mathbf{1}v^{\top},v\in\mathbb{R}^{n}$.
Thus, this completes our proof that $\left\{\mathbf{1}v^{\top}\mid v\in\mathbb{R}^{n}\right\}$ is the nullspace of $\phi$. ∎
@@
Proof B.23.
We start the proof by writing the arbitrary $H(x)\in\mathbf{P}_{[d-1]}^{n\times n}$ into its symmetric and skew-symmetric components $S(x)\in\mathbf{P}_{[d-1]}^{n\times n}$ and $\Omega(x)\in\mathbf{\Omega}_{[d-1]}^{n\times n}$ with $H(x)=S(x)+\Omega(x)$. For which, we can obtain by
@@
$=$ $S(x)x+x^{\top}S(x)x\mathbf{1}+\Omega(x)x.$
By Lemma 3.17, there exists a $B^{\prime}(x)\in\mathbf{\Omega}_{[d]}^{n\times n}$ such that $S(x)x+x^{\top}S(x)x\mathbf{1}=B^{\prime}(x)x$. Define $B(x):=B^{\prime}(x)+\Omega(x)$, we derive

### v2

@@ 3.2 The Reduction Technique
    There exists a $B\in\mathbf{\Omega}^{n\times n}_{[d]}$ such that $h(x)=B(x)x$, for all $x\in\mathbf{H}$.
@@
Proof 3.13.
First, we establish (16). Consider a linear change of variable $y=Tx$, for $T\in\mathbb{R}^{n\times n}$ nonsingular and $y_{1}:=x^{\top}\mathbf{1}$ (i.e., the first row of $T$ is $\mathbf{1}^{\top}$). Then, $x^{\top}g(x)$ can be written as a polynomial in $y$, which we denote by $q(y)$. Since $x^{\top}g(x)=0$ for all $x\in\mathbf{H}$, we have that $q(y)=0$ for all $y$ with $y_{1}=1$. It follows that $(1-y_{1})$ is a factor of $q(y)$ and hence, $(1-x^{\top}\mathbf{1})$ is a factor of $x^{\top}g(x)$.
@@
Proof 3.15.
Since $s(x)$ does not contain a constant or a linear term, we can write
@@
Proof 3.17.
The lemma follows directly from the construction of $s(x)$ and $h(x)$; indeed, for any $x\in\mathbb{R}^{n}$,
$x^{\top}g(x)=\left(1-x^{\top}\mathbf{1}\right)s(x)=\left(1-x^{\top}\mathbf{1}\right)x^{\top}H(x)x=x^{\top}h(x),$
where the first equality follows from Lemma 3.12, the second equality follows from Lemma 3.14, and the last equality follows from the definition of $h(x)$.
@@
where the third equality follows from the fact that $\mathbf{1}^{\top}x=1$ for $x\in\mathbf{H}$ and the last equality follows from the definition (17) of $h$. ∎
@@
As a consequence, if $H(x)$ is the payoff matrix (so $H(x)x$ is the payoff vector [park2019payoff]), then the associated replicator dynamics are given by
@@
where $g\in\mathbf{P}_{[2]}^{n}$, and $x^{\top}g(x)=\left(1-x_{1}-x_{2}-x_{3}\right)x_{1}^{2}+2x_{1}^{2}x_{2}-2x_{1}^{2}x_{2}=0$, for any $x\in\mathbf{H}$. We claim that there does not exist an $H\in\mathbf{P}_{[0]}^{n\times n}$ such that $g(x)$ can be written in the form of (19). Suppose, to the contrary, that such constant matrix $H$ exists; then, it follows that for any $i,j\in\{1,2,3\}$, the difference $g_{i}(x)-g_{j}(x)=(Hx)_{i}-(Hx)_{j}$ is a linear function in $x$, which contradicts the fact that $g_{2}(x)-g_{3}(x)=-x_{1}^{2}$ is quadratic. We thus conclude that no such constant $H$ exists for (19) to hold.
@@ 3.3 Proof of Theorem 1
The procedure of constructing the skew-symmetric $A(x)$ such that $g(x)=A(x)x$ is summarized in Algorithm 2, which has the same time complexity as Algorithm 1 (polynomial in $n$ and exponential in $d$).
@@ 8 Input $g^{\prime}(x)$ for Algorithm 1 to obtain $A^{\prime}(x)$
Output: Skew-symmetric payoff matrix $A(x)=A^{\prime}(x)+B(x)$
@@ 4 Conclusion and Outlook
A technical assumption we take is that the payoff vector $p(x)$ is polynomial. We believe that our result can be extended to the class of continuous payoff vectors. Since the simplex $\Delta^{n-1}$ is compact, by Stone-Weierstrass theorem [timan2014theory] any continuous function defined on the simplex can be approximated uniformly and arbitrarily well by polynomials. However, it remains unknown whether a convergent sequence of polynomial payoff vectors yields a convergent sequence of skew-symmetric, polynomial payoff matrices. We will address this issue in the future work. A related issue (and also a future direction) is about the robustness of Algorithm 2 against perturbation of the input $p(x)$, which matters when $p(x)$ is learned from data and, further, approximated by a polynomial.
@@ Appendix A Nullspace of map $\phi$
Let $\phi$ be given as in (2). Then, the nullspace of $\phi$, when restricted to $\mathbf{P}_{[0]}^{n\times n}$}, is $\left\{\mathbf{1}v^{\top}\mid v\in\mathbb{R}^{n}\right\}$.
Proof A.22.
@@
Proof of sufficiency. Let $H=\mathbf{1}v^{\top}$ with some $v\in\mathbb{R}^{n}$. For any $x\in\mathbb{R}^{n}$ with $x^{\top}\mathbf{1}=1$, we have
@@
$\phi(H)=Hx-\left(x^{\top}Hx\right)\mathbf{1}=\mathbf{1}\left(v^{\top}x\right)-\mathbf{1}\left(v^{\top}x\right)=0.$
@@
Proof of necessity. Let $H\in{\mathbf{P}_{[0]}^{n\times n}}$ be in the kernel of $\phi$, i.e., $\phi(H)(x)=0$ for all $x\in\mathbb{R}^{n}$ such that $\mathbf{1}^{\top}x=1$. Consider any $x$ in the relative interior of the simplex (i.e., $x_{i}>0$ for all $i$, $\mathbf{1}^{\top}x=1$). Since $\operatorname{diag}(x)$ is nonsingular, the condition $\phi(H)=0$ implies
$Hx=\left(x^{\top}Hx\right)\mathbf{1}.$ (23)
For any $i\neq j$, left-multiplying (23) by $(e_{i}-e_{j})^{\top}$ yields
$(e_{i}-e_{j})^{\top}Hx=\left(x^{\top}Hx\right)(e_{i}-e_{j})^{\top}\mathbf{1}=0,$ (24)
where the second equality follows from $(e_{i}-e_{j})^{\top}\mathbf{1}=0$. Now, consider any $y\in\mathbb{R}^{n}$ such that $\mathbf{1}^{\top}y=0$. Let $x_{0}$ be a point in the relative interior of the simplex. For a sufficiently small scalar $t$, $x_{0}+ty$ remains in the relative interior. By (24), we have
$(e_{i}-e_{j})^{\top}H(x_{0}+ty)=(e_{i}-e_{j})^{\top}Hx_{0}+t(e_{i}-e_{j})^{\top}Hy=0.$
Since $(e_{i}-e_{j})^{\top}Hx_{0}=0$, it follows that $(e_{i}-e_{j})^{\top}Hy=0$ and that $(e_{i}-e_{j})^{\top}H$ lies in the orthogonal complement of $\{y:\mathbf{1}^{\top}y=0\}$, which is $\operatorname{span}\{\mathbf{1}\}$. Hence there exists $c$ such that
$(e_{i}-e_{j})^{\top}H=c\mathbf{1}^{\top}.$
Right-multiplying by $x_{0}$ yields $0=c\mathbf{1}^{\top}x_{0}=c$, which implies that $c=0$. Therefore, $e_{i}^{\top}H=e_{j}^{\top}H$ for any $i\neq j$, i.e., $H=\mathbf{1}v^{\top}$ for some $v\in\mathbb{R}^{n}$. ∎
@@
Proof B.24.
We start the proof by writing $H(x)\in\mathbf{P}_{[d-1]}^{n\times n}$ into its symmetric and skew-symmetric components $S(x)\in\mathbf{P}_{[d-1]}^{n\times n}$ and $\Omega(x)\in\mathbf{\Omega}_{[d-1]}^{n\times n}$ with $H(x)=S(x)+\Omega(x)$. For which, we can obtain by
@@
$=$ $S(x)x-x^{\top}S(x)x\mathbf{1}+\Omega(x)x.$
where the second equation follows from $x^{\top}\Omega x=0$. By Lemma 3.18, there exists a $B^{\prime}(x)\in\mathbf{\Omega}_{[d]}^{n\times n}$ such that $S(x)x-x^{\top}S(x)x\mathbf{1}=B^{\prime}(x)x$. Define $B(x):=B^{\prime}(x)+\Omega(x)$, we derive